OpenStudy (anonymous):

3 sqrt108+ 3 sqrt1372 - 3 sqrt500

OpenStudy (anonymous):

@kaylarocks001

OpenStudy (anonymous):

ok so it is 3^108+3^1372-3^500 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well first you need to find out what each set equals.. 3^108=? 3^1372=? 3^3^500=?

OpenStudy (anonymous):

i dont have sqaure root on my calculater can you help me

OpenStudy (anonymous):

what are your choices?

OpenStudy (anonymous):

4.2 3sqrt4 15 3sqrt4 16 3sqrt4 5 3sqrt4

OpenStudy (anonymous):

find the general solution.. 1. y''-2y'-2y=0 2. y''''-8y''-9y=0 3. y''''+13y''+36y=0 4. y'''''-15y''''+84y'''-220y''-275y'=0 5. y'''''+12y''''+104y'''+408y''-1156y'=0

OpenStudy (anonymous):

@ash will you help

OpenStudy (anonymous):

@kaylarocks001 ..thts the ff qstion that i wud lyk 2 ans..2 u... find the general solution.. 1. y''-2y'-2y=0 2. y''''-8y''-9y=0 3. y''''+13y''+36y=0 4. y'''''-15y''''+84y'''-220y''-275y'=0 5. y'''''+12y''''+104y'''+408y''-1156y'=0